OpenStudy (anonymous):

a^2 + 12ab + 27b^2 Answer: (a+9b)(a+3b) I have to show my work for how to check the answer but im not sure how to do it.

OpenStudy (anonymous):

factroing trinomials btw

OpenStudy (ledah):

read this tutorial on how to factor. this will help better your understanding http://www.algebrahelp.com/lessons/factoring/trinomial/

OpenStudy (anonymous):

i know how to factor i just dont know how to check my answer for this one

jhonyy9 (jhonyy9):

so you need just multiplie term by term from parantheses the same like (a+b)(a+b)=a^2 +ab +ab +b^2 =a^2 +2ab +b^2 -so hope so much that is understandably sure right good luck bye